A sample of monoatomic hydrogen gas contains n atoms all in 3rd exited state.Find the minimum value of n so that on subsequent de-exitation all possible different energy photons are emitted?


Ans:4


 

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Use the energy equation

 E = -{k_e e^2 \over 2r_n } = - { (k_e e^2)^2 m_e \over 2\hbar^2 n^2} = {-13.6 \mathrm{eV} \over n^2}
 we will get n=4(min)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi